OpenStudy (anonymous):

Solve the equation 21=x-15

Vocaloid (vocaloid):

add 15 to both sides

OpenStudy (anonymous):

21=x-15 +15 +15 36 =×-30

OpenStudy (anonymous):

What do I do next is it correct

OpenStudy (anonymous):

Yes x=36

OpenStudy (anonymous):

So its correct

OpenStudy (anonymous):

yes

OpenStudy (kva1992):

the answer is sorta correct

OpenStudy (anonymous):

Its 36-15=21

OpenStudy (anonymous):

Can you show me how to do it then please

OpenStudy (anonymous):

not 30

OpenStudy (kva1992):

I mean the answer is correct but yeah its not 30 the 15 cancels

OpenStudy (anonymous):

Medal?

Vocaloid (vocaloid):

21=x-15 21 + 15 = x - 15 + 15 36 = x

OpenStudy (anonymous):

.................................

OpenStudy (kva1992):

there you go

OpenStudy (anonymous):

What about this one n-13=8

OpenStudy (kva1992):

same thing instead of an x its just a n treat it like you would treat a x

OpenStudy (anonymous):

So n-13=8 +13 +13 n- 26 =21

OpenStudy (kva1992):

nope n-13=8 +13 +13 n=21 -13 +13 =0

OpenStudy (kva1992):

dont forget to cancel

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

What about -44=w+32 -44 -44 W=-76

OpenStudy (kva1992):

yeah thats pretty much it

OpenStudy (kva1992):

try to keep the variable on its own side but you pretty much got it

OpenStudy (kva1992):

oh and also when you try to cancel you have to add the opposite from one side so if its -44 you add 44 to cancel not subtract 44